OpenStudy (anonymous):

Balance the following equations: a.) Al(OH)3 + H2SO4 ---> Al2(SO4)3 + H2O b.) S8 + O2 ---> SO3 c.) SiCl4 + H2O ---> H4SiO4 + HCl d.) Na2CO3 + HCl ---> NaCl + H2O +CO2 e.) KOH +H3PO4 ---> K3PO4 + H2O

OpenStudy (anonymous):

2 Al(OH)3 + 3 H2SO4 = Al2(SO4)3 + 6 H2O S8 + 12 O2 = 8 SO3 SiCl4 + 4 H2O = H4SiO4 + 4 HCl Na2CO3 + 2 HCl = 2 NaCl + H2O + CO2 3 KOH + H3PO4 = K3PO4 + 3 H2O